OpenStudy (anonymous):

Integrate x^5 * sqrt(1 - 4x^2) No Trig substitution. What should my first few steps be?

OpenStudy (chihiroasleaf):

is this your question? \[\int\limits x ^{5}\sqrt{1-4x ^{2}} dx\] use partial intergration

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

Should the u be x^5 or the sqrt section, or will both work equally well?

OpenStudy (chihiroasleaf):

yes. lt u = x^ 5 and dv= sqrt(1 - 4x^2) dx you'll do several integration by parts to find the answer...

OpenStudy (raden):

i think that problem can be using u-subt method look here :

OpenStudy (raden):